An inductance of L Henrys and a resistance of 10 ohms are connected in series with e.m.f. of 100 volts. If the current is initially zero, and is equal to 9 amperes after 1 second , find L and find the current after 0.5 second

1
Expert's answer
2022-03-07T15:51:02-0500

The instantenous emf equation is given as;

"L\frac{di}{dt}+Ri=V"

"\frac{di}{dt}=\frac{V-Ri}{L}"

"\frac{di}{dt}=\frac{R}{L}(\frac{V}{R}-i)"

"\frac{di}{(\frac{V}{R}-i)}=\frac{R}{L}dt"

Integrating on both sides

"\int_0^i \frac{di}{(\frac{V}{R}-i)}=\int_0^t \frac{R}{L}dt"

"\int_0^i \frac{di}{(\frac{i-V}{R})}=-\frac{R}{L}\int_0^tdt"

"ln\frac{(i-\frac{V}{R})}{-\frac{V}{R}}=-\frac{R}{L}t"

"\frac{(i-\frac{V}{R})}{-\frac{V}{R}}=e^{-\frac{R}{L}t}"

"i={\frac{V}{R}}(1-e^{-\frac{R}{L}t})"

When t=1, i(1)=9

"9={\frac{100}{10}}(1-e^{-\frac{10}{L}1})"

"\frac{9}{10}=1-e^{\frac{-10}{L}}"

"e^\frac{-10}{L}=0.1"

"\frac{-10}{L}=ln(0.1)"

ln(0.1)=-2.302

"L=\frac{10}{2.302}"

L=4.443H

Current after 0.5s is;

"i(0.5)={\frac{100}{10}}(1-e^{-\frac{10}{4.344}*0.5})"

=6.837A